Determine the solubilility of oxygen gas in water at 25 degrees celcius when the partial presssure of oxygen is 459.6mmHg. KH=3.40x10-2M/atm

I know the equation is S=KHP

I have the answer as 0.0206M but I cant get to this answer.
Answer #1

KH=3.40x10-2M/atm

P     = 459.6mmHg

        = 459.6/760   = 0.6047atm

S=KHP

= 3.4*10^-2 M/atm* 0.6047 atm

   = 0.0206M

The solubility of oxygen = 0.0206M >>>>answer